# SEE-DPO: Self Entropy Enhanced Direct Preference Optimization

### v1

@@ 4 Methodology
In this section, we prove the equivalence of D3PO, SPO, and DiffusionDPO. We model the diffusion process as an MDP, we use the notations from [39]:
@@
Where T is the maximum number of denoising timesteps and $\delta_{x}$ represents a Dirac delta distribution. We want to optimize Eq. (9), we can re-write it as:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{b}}\Bigg{[}\log\sigma\Bigg{(}$
$\beta\log\prod_{i}\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}$
@@
$-\beta\log\prod_{i}\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg{)}\Bigg{]}$ (10)
We don’t write the conditioning on c explicitly in Eq. (10) for conciseness. This can be further written as:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{b}}\Bigg{[}\log\sigma\Bigg{(}$
$\beta\sum_{t=1}^{N}\log\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}$
$-\beta\sum_{t=1}^{M}\log\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg{)}\Bigg{]}$ (11)
Where $N,M$ are the number of timesteps in the winning and losing sample generation respectively.
We assume that $N=M=T$, and using [39], we can further write the objective as:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{b}}\Bigg{[}\log\sigma\Bigg{(}$
$\beta T*\frac{1}{T}*\sum_{t=1}^{T}\log\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}$
$-\beta T*\frac{1}{T}*\sum_{t=0}^{T}\log\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg{)}\Bigg{]}$ (12)
@@
We can define t as a random variable which is sample from a uniform distribution with support from 0 to T. Using this and jenson shannon inequality we arrive at:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{b}}\Bigg{[}\log\sigma\Bigg{(}$
@@
$\beta T*\mathbb{E}_{t\sim\mathcal{U}(0,T)}\Bigg{[}\log\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}$
$-\log\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg{]}\Bigg{)}\Bigg{]}$ (13)
$L_{\text{DPO}}(\theta)\leq-\mathbb{E}_{c,x_{0}^{w},x_{0}^{b},t\sim\mathcal{U}(0,T)}\Bigg{[}\log\sigma\Bigg{(}$
$\beta T\log\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}$
$-\beta T\log\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg{)}\Bigg{]}$ (14)
Figure 2: Accuracy vs Steps A: Step-Wise PickScore [22], B: Aesthetic Score[35], C: Pick Score [18], demonstrate faster convergence, improved training stability and robustness to reward hacking.
@@
This objective is similar to Diffusion-DPO, if instead we consider optimization at each step then the objective translates to SPO and D3PO’s objective.
Now we introduce our self-entropy regularization into this framework. We start with D3PO’s formulation:
       DPOK 20.583 0.260 4.705 0.498 0.099 0.270
       DDPO 20.569 0.260 4.683 0.498 0.090 0.270
       D3PO 20.306 0.227 5.050 0.470 -0.209 0.261
       DiffusionDPO 20.791 0.262 4.711 0.486 0.139 0.263
       SPO 20.919 0.266 4.702 0.472 0.191 0.249
       SEE-D3PO 20.590 0.247 4.851 0.491 -0.003 0.271
       SEE-DiffusionDPO 20.854 0.270 4.714 0.491 0.225 0.265
       SEE-SPO 21.256 0.283 4.673 0.509 0.590 0.273
Table 1: All numbers are reported on the validation-unique split of Pick-a-Pic V1 dataset. We evaluate on CLIP Score [13], BLIP Score [20], Aesthetic Score [35], Pick Score [18], HPSv2 Score [36]
$\max_{\pi}\mathbb{E}_{s\sim d^{\pi},a\sim\pi(\cdot|s)}\Big{[}Q^{*}(s,a)\Big{]}-$
$\beta\Bigg{(}D_{\text{KL}}\Big{[}\pi(a|s)\|\pi_{\text{ref}}(a|s)\Big{]}+\gamma\pi(a|s)\log{\frac{\pi(a|s)}{\pi_{\text{ref}}(a|s)}}\Bigg{)}$ (15)
Given the objective of Eq. (15), the action-value function can be written in term of the optimal policy $\pi^{*}(a|s)$ as:
$Q^{*}(s,a)=\beta\left(\log{\frac{\pi^{*}(a|s)}{\pi_{\text{ref}}(a|s)}}+\gamma\log{\pi^{*}(a|s)}\right)$ (16)
Plugging Eq. (16) in the Bardley-Terry Model and optimizing the log-likelihood we get the following objective function:
@@
$\mathcal{L}(\theta)=-\mathbb{E}_{(s_{k},\sigma_{w},\sigma_{t})}\Bigg{[}\log\rho\Bigg{(}\beta(\gamma+1)\quad\cdot$
$\hskip 28.45274pt\Bigg{(}\log\frac{\pi_{\theta}(a_{k}^{w}|s_{k}^{w})}{\pi_{\text{ref}}(a_{k}^{w}|s_{k}^{w})^{\frac{1}{\gamma+1}}}-\log\frac{\pi_{\theta}(a_{k}^{t}|s_{k}^{t})}{\pi_{\text{ref}}(a_{k}^{t}|s_{k}^{t})^{\frac{1}{\gamma+1}}}\Bigg{)}\Bigg{)}\Bigg{]}$ (17)
@@
From Eq. (17) we can see that the final effect of adding the self-entropy term was to flatten out the reference distribution which in return forces the policy to be more explorative. We can also write this in terms of the noise objective similar to Diffusion-DPO:
@@
We could start from the formulation of Diffusion-DPO to get an alternate version of our objective which is different only by a constant, we start by optimizing the below objective:
`$\max_{p_{\theta}}\mathbb{E}_{c\sim\mathcal{D}_{c},x_{0:T}\sim p_{\theta}(x_{0:T}\mid c)}[r(c,x_{0})$
$\hskip 56.9055pt-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0:T}\mid c)\parallel p_{\text{ref}}(x_{0:T}\mid c)]$ (18)

### v2

@@ 4 Methodology
In this section, we prove the equivalence of D3PO, SPO, and Diffusion-DPO. We model the diffusion process as an MDP, we borrow the notations from Yang et al. (2024):
@@
Where $T$ is the maximum number of denoising timesteps and $\delta_{x}$ represents a Dirac delta distribution. We want to optimize Eq. 9, we can re-write it as:
@@
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l}}\Bigg[\log\sigma\Bigg(\beta\log\prod_{i}\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}-\beta\log\prod_{i}\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg)\Bigg]$ (13)
We assume that the number of genration timestep ($T$) is same for both winning and lossing sample, by using the property $\log{x1\cdot x2}=\log{x1}+\log{x2}$ $\forall x1,x2\in R^{+}$ we simplify 13. We don’t write the conditioning on $c$ explicitly in Eq. 13 for conciseness:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l}}\Bigg[\log\sigma\Bigg(\beta T*\frac{1}{T}*\sum_{t=1}^{T}\log\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}-\beta T*\frac{1}{T}*\sum_{t=0}^{T}\log\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg)\Bigg]$ (14)
@@
We can define $t$ as a random variable which is sampled from a uniform distribution with support from 0 to $T$. Using this and Jenson Shannon inequality we arrive at:
@@
$L_{\text{DPO}}(\theta)\leq-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l},t\sim\mathcal{U}(0,T)}\Bigg[\log\sigma\Bigg(\beta T\log\frac{p_{\theta}(x_{t-1}^{w}|x_{t}^{w})}{p_{\text{ref}}(x_{t-1}^{w}|x_{t}^{w})}-\beta T\log\frac{p_{\theta}(x_{t-1}^{l}|x_{t}^{l})}{p_{\text{ref}}(x_{t-1}^{l}|x_{t}^{l})}\Bigg)\Bigg]$ (15)
@@
This objective is similar to Diffusion-DPO, if instead we consider optimization at each step then the objective translates to SPO and D3PO’s objective. Now we introduce our self-entropy regularization into this framework. We start with D3PO’s formulation:
$\max_{\pi}\mathbb{E}_{s\sim d^{\pi},a\sim\pi(\cdot|s)}\Big[Q^{*}(s,a)\Big]$ $-\beta\Bigg(D_{\text{KL}}\Big[\pi(a|s)\|\pi_{\text{ref}}(a|s)\Big]+\gamma\pi(a|s)\log{\pi(a|s)}\Bigg)$
$\max_{\pi}\mathbb{E}_{s\sim d^{\pi},a\sim\pi(\cdot|s)}\Big[Q^{*}(s,a)\Big]$ $-\beta\Bigg(\pi(a|s)\log{\frac{\pi(a|s)}{\pi_{\text{ref}}(a|s)}}+\gamma\pi(a|s)\log{\pi(a|s)}\Bigg)$
$\max_{\pi}\mathbb{E}_{s\sim d^{\pi},a\sim\pi(\cdot|s)}\Big[Q^{*}(s,a)\Big]$ $-\beta\Bigg((\gamma+1)\pi(a|s)\log{\frac{\pi(a|s)}{{\pi_{\text{ref}}(a|s)}^{\frac{1}{\gamma+1}}}}\Bigg)$
$\max_{\pi}\mathbb{E}_{s\sim d^{\pi},a\sim\pi(\cdot|s)}\Big[Q^{*}(s,a)\Big]$ $-\beta(\gamma+1)\Bigg(D_{\text{KL}}\Big[\pi(a|s)\|{\pi_{\text{ref}}(a|s)}^{\frac{1}{\gamma+1}}\Big]\Bigg)$ (16)
We can see that the only effect of adding our self entropy term is on $\pi_{ref}$, particularly we can simplify Eq. 16 using the usual DPO derivation Rafailov et al. (2024b) to show that adding the self entropy is equivalent to raising $\pi_{ref}$ by $1+\gamma$. Hence we can use the same procedure as D3PO to arrive at our update objective:
@@
$L(\theta)=-\mathbb{E}_{(s_{k},\sigma_{w},\sigma_{t})}\Bigg[\log\rho\Bigg(\beta(\gamma+1)\cdot\Bigg(\log\frac{\pi_{\theta}(a_{k}^{w}|s_{k}^{w})}{\pi_{\text{ref}}(a_{k}^{w}|s_{k}^{w})^{\frac{1}{\gamma+1}}}-\log\frac{\pi_{\theta}(a_{k}^{t}|s_{k}^{t})}{\pi_{\text{ref}}(a_{k}^{t}|s_{k}^{t})^{\frac{1}{\gamma+1}}}\Bigg)\Bigg)\Bigg]$ (17)
@@
Since SPO follows the same formulation as D3PO Eq. 17 can also be applied to train SPO. From Eq. (17) we can see that for $\gamma>0$ the final effect of adding the self-entropy term was to flatten out the reference distribution which in return forces the policy to be more explorative (appendix C.1), we can also write this in terms of the noise objective, similar to Diffusion-DPO:
@@
We could also start from the formulation of Diffusion-DPO to get an alternate version of our objective which is different only by a constant, we follow the same steps as of Diffusion-DPO to arrive at our final objective the exact derivation can be found in the appendix C.2:
We can see that to a constant factor this is similar to one derived directly from Eq. 17. Intuitively, in this formulation, the loss associated with the reference model is reduced when $\gamma>0$. Therefore the resulting diffusion model relies less on the reference model. This allows the trained model to explore the low probability space of the reference model and thus increases the diversity of the trained model.
5 Experiments
Figure 2: Accuracy vs Steps A: Step-Wise PickScore Liang et al. (2024), B: Aesthetic ScoreWang et al. (2022), C: PickScore Kirstain et al. (2023), demonstrate faster convergence, improved training stability and robustness to reward hacking.
DiversityPrompt AlignmentVisual AppealGeneral Preference32.532.527.727.732.532.532.332.329.129.140.540.523.223.225.125.138.438.431.831.844.344.342.642.6SEE-SPO WinDrawSPO WinPrompt AlignmentVisual AppealGeneral PreferenceDiversity25.525.527.327.332.732.733.433.441.741.740.240.221.721.723.423.432.832.832.532.545.645.643.243.2SEE-DiffusionDPO WinDrawDiffusionDPO Win02020404060608080100100DiversityPrompt AlignmentVisual AppealGeneral Preference30.130.126.126.127.927.930.630.627.727.739.839.820.620.618.218.242.242.234.134.151.551.551.251.2SEE-D3PO WinDrawD3PO Win
Figure 3: User study. We compare our best models with counterparts fine-tuned using the PickScore reward model Kirstain et al. (2023). Following SPO Liang et al. (2024), we randomly select 300 prompts from PartiPrompts Yu et al. (2022) and HPS Wu et al. (2023) in a 1:2 ratio. Participants are shown a reference image alongside three generated images from the same prompt but with different initial latents. They evaluate image quality and diversity relative to the reference to estimate the diversity metric. For other metrics, users assess single images.
    D3PO 20.306 0.227 5.050 0.470 -0.209 0.261
    Diffusion-DPO 20.791 0.262 4.711 0.486 0.139 0.263
    SPO 20.919 0.266 4.702 0.472 0.191 0.249
    SEE-D3PO 20.590 0.247 4.851 0.491 -0.003 0.271
    SEE-DiffusionDPO 21.014 0.273 4.721 0.497 0.333 0.266
    SEE-SPO 21.295 0.287 4.679 0.510 0.585 0.275
Table 1: Quality Metrics.All numbers are reported on the validation-unique split of Pick-a-Pic V1 dataset. We evaluate on CLIP Score Hessel et al. (2022), BLIP Score Li et al. (2022), Aesthetic Score Wang et al. (2022), PickScore Kirstain et al. (2023), HPSv2 Score Wu et al. (2023)
    StableDiffusion-1.5 0.3745 8.6943 0.2100 7.4627 7.4620
    Diffusion-DPO 0.3736 8.7047 0.2406 7.4954 7.4945
    SEE-DiffusionDPO 0.4299 8.1743 0.2056 7.6517 7.6587
    D3PO 0.2455 12.5599 0.4190 7.0853 7.0911
    SEE-D3PO 0.3052 10.5240 0.3183 7.3650 7.3692
    SPO 0.4446 7.1218 0.1666 7.5674 7.5635
    SEE-SPO 0.4383 7.0818 0.1575 7.7120 7.7101
Table 2: Diversity Metrics.All values are reported on the validation-unique split of the Pick-a-Pic-V1 dataset Kirstain et al. (2023). For each prompt, we first sample a base image and then sample 10 different images to get all the diversity metrics. Evaluation metrics include RMSE, PSNR, SSIM, and entropy energy scores (E1 and E2) Sun et al. (2024).
Experimental Setting: For a fair comparison, we used the official implementations of D3PO, Diffusion-DPO, and SPO with their default parameters. We trained these models using our proposed regularized loss function, as described in the Methodology section. When applying our method, we treated only $\gamma$ and $\beta$ as hyperparameters while keeping all other settings at their default values to ensure a fair evaluation. During inference, we set the guidance scale to 7.5 for consistency across models.
For training, we used 4,000 prompts from the Pick-a-Pic-V1 dataset Kirstain et al. (2023) for SPO and D3PO, following the dataset provided in SPO’s official GitHub repository. For Diffusion-DPO, we used 800,000 prompts from the same dataset. Each model was trained using the same setup and data splits as specified in their original implementations. To establish a fair basis for comparison, we selected Stable-Diffusion v1.5 as the benchmark model.
Since our training approach is online, we relied on automatic preference generators. Specifically, we used the time-conditional PickScore Liang et al. (2024) as a proxy for human preference in SPO, human annotations from the Pick-a-Pic-V1 dataset for Diffusion-DPO, and PickScore Kirstain et al. (2023) for D3PO. Notably, we did not use online training for Diffusion-DPO to remain consistent with its original implementation.
Evaluation: To assess the quality of generated images, we utilize automatic evaluation metrics, including PickScore Kirstain et al. (2023), HPS Wu et al. (2023), Aesthetic Score Wang et al. (2022), CLIP Hessel et al. (2022), BLIP Li et al. (2022), and ImageReward Xu et al. (2023). We report results on the validation_unique split of the Pick-a-Pic V1 dataset, which contains 500 prompts, as shown in Tables 1, 2, 4 and 5.
For a comprehensive assessment of generation diversity, we employ five additional metrics. Root Mean Square Error (RMSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index Measure (SSIM) are commonly used to evaluate image similarity; here, we leverage them to quantify diversity in generated images. Additionally, we compute image entropy, including both 1D and 2D entropy, to measure the information content within images. Higher entropy values indicate greater diversity and richness in image content.
To evaluate diversity, we generate 10 images per prompt from the validation_unique set and compute the aforementioned diversity metrics Sun et al. (2024). We do not rely on embedding-based similarity, as images generated from the same prompt, despite visual differences, often cluster closely in embedding space. An effectively aligned model should achieve high reward scores while producing diverse outputs. Our results demonstrate that the regularized versions of D3PO, SPO, and Diffusion-DPO achieve higher mean rewards across multiple metrics while improving diversity scores, highlighting enhanced performance in both diversity and image quality compared to baseline methods.
Additionally, we conduct a user study similar to Liang et al. (2024). We recruit 10 participants to evaluate images generated by different models based on 300 prompts sampled from PartiPrompts and HPSv2 in a 1:2 ratio. Participants assess four aspects: Prompt Alignment, Visual Appeal, General Preference, and Diversity. To evaluate diversity, we provide each participant with a reference image generated by a model alongside three additional images generated from different latent codes. Participants rate both their general preference and the diversity of the additional images based on their differences from the reference image.
Model $\gamma$ $\beta$
D3PO 5 0.01
Diffusion-DPO 3 4
SPO 3 0.1
5.1 Results
Our models, trained using the proposed objective, outperform or are comparable to baseline methods across all quality metrics, as shown in Table 2. Specifically, SEE-D3PO improves PickScore by 1.41%, HPS by 8.8%, and ImageReward by 98.5%. SEE-DiffusionDPO yields improvements of 1.08% in PickScore, 4.2% in HPS, and 139.5% in ImageReward. Lastly, SEE-SPO shows gains of 1.8% in SPO, 7.9% in HPS, and 206.3% in ImageReward.
In terms of image diversity metrics, all our models outperform their respective counterparts across all six metrics. This trend is also reflected in the user study [Fig: 3], where our methods achieve at least a 6% improvement in generation diversity. Furthermore, our approach enhances all other evaluation aspects, though we observe minimal gains in the Prompt Alignment task. Example samples are provided in the appendix and Fig: 4.
5.2 Ablations
Figure 5: Ablation Study for $\gamma$ & $\beta$: This study investigates the impact of $\gamma$ and $\beta$ on image quality, specifically using PickScore as the metric. In the first ablation, we examine the effect of varying $\gamma$ while keeping all other parameters fixed, with $\beta$ set to 0.1. In the second ablation, we explore the influence of different values of $\beta$ for $\gamma=3$ settings.
In our experiments, we explored the effect of different values of $\gamma$ on both training stability and overall model performance. As shown in Fig: 2, when $\gamma=0$, representing the baseline or vanilla model, we observe an initial steep drop in reward values, followed by a sudden and rapid increase during training phases of all the models SPO and D3PO training phases. Examining the generated images in this region, we note a gradual decline in quality corresponding to the initial drop in scores, followed by further degradation but the reward values continue to rise. This pattern signals that the model begins to overfit prematurely, leading to the generation of repetitive and noisy images—a phenomenon we define as reward hacking. In this scenario, the model exploits reward dynamics by producing images that maximize the reward function but lead to distorted images as shown in Fig: 1. Through our experimentation, we observed two distinct types of overfitting. In the first, seen in D3PO, the reward values continue to increase despite a progressive decline in image quality, indicating that the model is optimizing for reward at the expense of meaningful outputs. In the second, observed in SPO, the reward model assigns consistently low scores, and the model fails to improve beyond this point, becoming stuck in a suboptimal state. Notably, this trend is absent in the Diffusion-DPO setup, which utilizes the extensive Pick-a-Pic-V1 dataset containing over 800,000 images, likely providing more comprehensive training data that mitigates early overfitting. For diffusion DPO we see that the model with $\gamma=-0.5$ initially increases in accuracy and then decreases. Across all configurations, our model achieves higher scores faster, with more consistent stability in training and fewer abrupt changes in reward values.
As training progresses, our model not only stabilizes but also consistently reaches a higher reward range. We attribute this improvement to the implicit flattening of the reference distribution that higher $\gamma$ values introduce. This flattening effect broadens the policy distribution, enabling more exploratory behavior and decreasing the likelihood of overfitting to repetitive patterns.
Further, we observe empirically poorer results compared to the base model when $\gamma=-0.5$, as shown in Fig: 3. D3PO demonstrates early overfitting tendencies, resulting in sharp reward increases followed by rapid declines. Meanwhile, Diffusion-DPO consistently underperforms relative to other models, as the negative $\gamma$ value restricts exploratory actions, yielding repetitive outputs that fail to capture meaningful image variations. In the SPO pipeline, overfitting occurs somewhat later than in the vanilla model. This difference can be attributed to SPO’s update policy, which only updates the model when it generates images with reward differences that exceed a predetermined threshold. With a negative $\gamma$, the generation distribution narrows, limiting the diversity of generated images and, consequently, the frequency of model updates. Further evidence of these observations can be seen by examining the rewards vs. steps graph for SPO $\gamma=-0.5$, where the reward values remain nearly constant throughout most of the training duration before ultimately showing signs of overfitting. This steady reward trend reinforces the observation that the model fails to explore diverse solutions, hindering its ability to generalize and adapt, particularly under negative $\gamma$ settings.
The optimal values for $\beta$ and $\gamma$ are shown in Table 3. Notably, the optimal $\beta$ for Diffusion-DPO is two orders of magnitude larger than that of D3PO, which is expected since the $\beta$ for Diffusion-DPO is multiplied by $T$, a term of order 2. In the case of SPO, we update at each timestep, which is of order 1, so we also expect $\beta$ to increase by one order to prevent significant deviation from the reference distribution.
We also conducted a deeper ablation study using SPO, exploring various values of $\beta$ and $\gamma$. In Fig: 5, the left image shows the results of keeping $\beta$ fixed at 0.1 while varying $\gamma$. We observe that for $\gamma>1$, all experiments achieve high PickScore, whereas for $\gamma<0$, the performance significantly drops. As noted earlier, $\gamma<0$ leads to reduced diversity and poor training, with results comparable to the vanilla Stable-Diffusion 1.5 model. In the right image, we fix $\gamma=3$ and vary $\beta$ while keeping other hyperparameters constant. Initially, for low values of $\beta$, the performance is high, but it declines steeply as $\beta$ increases. This decline may be due to the model’s increasing reliance on the initial distribution as the regularization strength $\beta$ rises, limiting exploration and slowing training, which results in lower scores.
6 Conclusion
In this paper, we introduced Self-Entropy Enhanced Direct Preference Optimization (SEE-DPO) for fine-tuning diffusion models. We showed existing approaches that optimize Eq. 5 often suffer from overfitting and distribution shift issues. To address these challenges, we introduced a self-entropy regularization term into the objective function, transforming it into the final optimization objective outlined in Eq. 17. Our empirical results demonstrate that the proposed SEE-DPO, SEE-DiffusionDPO, and SEE-DPO method significantly outperforms their respective counterparts, effectively mitigating the reward hacking problem commonly associated with RLHF approaches and resulting in robust training. Additionally, our method ensures diversity of image generation across the latent space, highlighting its robustness and effectiveness in maintaining high-quality outputs.
Our approach was straightforward to implement within the current framework, specifically DPO with KL-regularized reward optimization. However, extending our method to incorporate other forms of regularization, such as Hellinger distance, may introduce additional mathematical complexity. We leave the exploration of alternative regularizations and their impact on performance to future work. Furthermore, although our diversity metrics (RMSE and PSNR) showed minimal improvement with SPO, the user study reported a higher diversity. This discrepancy highlights the need for a diversity metric that better aligns with human judgment, which we plan to investigate in future work.
7 Impact Statement
Generative models for media, such as image generation, present both significant opportunities and challenges. On one hand, these models enable a wide range of creative applications and, with advancements that reduce training and inference costs, have the potential to make this technology more accessible and encourage broader exploration. On the other hand, they also pose risks by making it easier to create and disseminate manipulated data, which can contribute to the spread of misinformation and harmful content, including the prevalent issue of deepfakes. Additionally, generative models may inadvertently expose sensitive or personal information from their training data, raising concerns about privacy, particularly when such data is collected without consent. The potential for this issue in diffusion models, specifically in image generation, remains an area requiring further investigation. Moreover, deep learning systems, including diffusion models, can perpetuate or amplify biases inherent in their training data. Our methods explores increasing diversity of generation which could positively impact the bias-corretion problem. For a more detailed discussion of the ethical considerations of deep generative models we refer readers to Hagendorff (2024).
Appendix A Appendix
This supplementary document provides additional details to complement the main paper:
  • Additional experiment results
  • Additional qualitative results
  • Proofs
Appendix B Additional Experiment Results
In this section, we present additional evidence of the statistical significance of our results, as shown in Tables 4 and 5. All experiments were conducted over five randomly chosen seeds to ensure robustness. Furthermore, for Fig: 6, we generated five images per prompt using different latent variables while keeping the prompts consistent.
Our model consistently outperforms its counterparts, producing images with significantly finer details, thereby demonstrating its robustness and effectiveness in generating high-quality outputs.
Among the evaluated visual metrics, we observe that SSIM consistently aligns with the visual improvements, we attribute this to the fact that it effectively measures fine-grained structural similarities. In contrast, RMSE and PSNR, being coarse-grained metrics, failed to account for the nuanced details in the generated images. These results highlight the importance of using metrics like SSIM for a more comprehensive evaluation of fine-grained image generation tasks.
    D3PO $20.306\pm 0.016$ $0.227\pm 0.001$ $5.050\pm 0.005$ $0.470\pm 0.001$ $0.209\pm 0.038$ $0.261\pm 0.002$
    Diffusion-DPO $20.791\pm 0.021$ $0.262\pm 0.001$ $4.711\pm 0.010$ $0.486\pm 0.002$ $0.139\pm 0.034$ $0.263\pm 0.001$
    SPO $20.919\pm 0.024$ $0.266\pm 0.002$ $4.702\pm 0.019$ $0.472\pm 0.002$ $0.191\pm 0.028$ $0.249\pm 0.001$
    SEE-D3PO $20.590\pm 0.017$ $0.247\pm 0.001$ $4.851\pm 0.016$ $491\pm 0.001 0.0$$003\pm 0.018$ $0.271\pm 0.002$
    SEE-DiffusionDPO $21.014\pm 0.032$ $0.273\pm 0.003$ $4.721\pm 0.015$ $0.497\pm 0.002$ $0.333\pm 0.052$ $0.266\pm 0.001$
    SEE-SPO $21.295\pm 0.041$ $0.283\pm 0.001$ $4.679\pm 0.010$ $0.510\pm 0.001$ $0.585\pm 0.028$ $0.275\pm 0.001$
Table 4: Results are averaged over 5 random seeds, with mean and standard deviation reported.All numbers are reported on the validation-unique split of Pick-a-Pic V1 dataset. We evaluate on CLIP Score, BLIP Score , Aesthetic Score , PickScore , HPSv2 Score
    D3PO $0.2455\pm 0.0006$ $12.5599\pm 0.0248$ $0.4190\pm 0.0014$ $7.0853\pm 0.0075$ $7.0911\pm 0.0074$
    Diffusion-DPO $0.3736\pm 0.0017$ $8.7047\pm 0.0401$ $0.2406\pm 0.0013$ $7.4954\pm 0.0071$ $7.4945\pm 0.0070$
    SPO $0.4446\pm 0.0008$ $7.1218\pm 0.0154$ $0.1666\pm 0.0015$ $7.5674\pm 0.0118$ $7.5635\pm 0.0118$
    SEE-DiffusionDPO $0.4299\pm 0.0005$ $8.1743\pm 0.0095$ $0.2056\pm 0.0025$ $7.6517\pm 0.0059$ $7.6587\pm 0.0060$
    SEE-D3PO $0.3052\pm 0.0003$ $10.5240\pm 0.0100$ $0.3183\pm 0.0021$ $7.3650\pm 0.0123$ $7.3692\pm 0.0124$
    SEE-SPO $0.4383\pm 0.0021$ $7.0818\pm 0.0351$ $0.1575\pm 0.0018$ $7120\pm 0.0072 7.7$$7101\pm 0.0074$
Table 5: Results are averaged over 5 random seeds, with mean and standard deviation reported.All values are reported on the validation-unique split of the Pick-a-Pic-V1 dataset. For each prompt, we first sample a base image and then sample 10 different images to get all the diversity metrics. Evaluation metrics include RMSE, PSNR, SSIM, and entropy energy scores (E1 and E2).
Figure 6: Images are generated using the same prompts across 5 different latents, demonstrating that our models consistently outperform their counterparts across all cases.
Appendix C Proofs
C.1 Effects of $\gamma$
The parameter $\gamma$ serves to flatten the reference distribution. As seen in Eq. 16, the objective includes a KL divergence term that penalizes deviation from the reference policy, which in turn can restrict the training policy from exploring potentially higher-reward regions. By flattening the reference distribution via a higher $\gamma$, the training policy is allowed greater flexibility to explore a wider range of actions. This effect is illustrated in Fig. 7, where the high-reward region is initially assigned very low probability under the standard reference policy, making it difficult for the training policy to reach it—resulting in slower learning. In contrast, with a larger $\gamma$, the reference distribution becomes more uniform, assigning greater probability mass to the high-reward region. This facilitates more effective exploration and accelerates training.
Figure 7: A toy example illustrating how incorporating $\gamma$ encourages greater exploration, potentially guiding the model toward regions with higher rewards.
C.2 Diffusion-DPO equivalence
We could also start from the formulation of Diffusion-DPO to get an alternate version of our objective which is different only by a constant, we follow the same steps as of Diffusion-DPO to arrive at our final objective:
$\max_{p_{\theta}}\mathbb{E}_{c\sim\mathcal{D}_{c},x_{0:T}\sim p_{\theta}(x_{0:T}\mid c)}[r(c,x_{0})-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0:T}\mid c)\parallel p_{\text{ref}}(x_{0:T}\mid c)]-\beta\gamma p_{\theta}(x_{0:T}\mid c)\log{p_{\theta}(x_{0:T}\mid c)}$ (18)